For each equation find teh x - intercept
y = x^2 + x - 6

Answer by stanbon(75887)

For each equation find teh x - intercept
y = x^2 + x - 6
-------------
To find the x-intercept let y = zero:
x^2+x-6 = 0
Factor to get:
(x+3)(x-2) = 0
Since the product is zero one of the factors must be zero:
x+3=0 or x-2=0
x=-3 or x=2
======================
Cheers,
Stan H.